5 years ago Jane was 5 times as old as her son. 5 years hence her age will be 8 less than 3 times the corresponding age of her son.Find their ages
Answer by ankor@dixie-net.com(22740)

let j = J's present age
Let s = son's present age
:
Write an equation for each statement, simplify each equation:
:
"5 years ago Jane was 5 times as old as her son."
j - 5 = 5(s-5)
j - 5 = 5s - 25
j = 5s - 25 + 5
j = 5s - 20
:
"5 years hence her age will be 8 less than 3 times the corresponding age of her son."
j + 5 = 3(s+5) - 8
j + 5 = 3s + 15 - 8
j + 5 = 3s + 7
j = 3s + 7 - 5
j = 3s + 2
:
From the 1st simplified equation, replace j with (5s-20) in the above eq
5s - 20 = 3s + 2
5s - 3s = 2 + 20
2s = 22
s = 22/2
s = 11 yrs is the son's present age
:
Find J's age
j = 3(11) + 2
j = 33 + 2
j = 35 is Jane's present age
:
:
Confirm this using the 1st statement:
"5 years ago Jane was 5 times as old as her son. "
35 - 5 = 5(11 - 5)
30 = 5(6)